OpenStudy (anonymous):

the product of two consecutive numbers is 42. what is the larger number?

OpenStudy (anonymous):

You know for a number n, its 'consecutive' number is n+1. This means that: \[n(n+1) = 42 \implies n^2 + n - 42 = 0 \implies (n_2-7)(n_1-6)=0 \implies n_{1,2} = 6,7\]

OpenStudy (anonymous):

The last part is:\[\implies n_{1,2} = 6,7\]

OpenStudy (anonymous):

thank you!!